OpenStudy (anonymous):

two sides of a triangle have lengths of 14 and 19 inches. Their included angle measures 82 degrees. Find the area of the triangle.

OpenStudy (nali):

use the Pythagorean theory a^2 + b^2 = c^2

OpenStudy (anonymous):

how are you supposed to use pythagorean? It's not a right triangle

OpenStudy (nali):

can u plz draw a picture of the problem

OpenStudy (anonymous):

 To be safe don't assume it's a right triangle

OpenStudy (nali):

in order to find the third side use the formula and let a= 14 and b = 19 \[ c^{2}= a^{2}+ b^{2} - 2ab\] cos C = \[14^{2}+ 19^{2} - 2(14)(19) cos 82 = 196 + 361 - (532) (.9497) = 51.7596 the answer is about 52

OpenStudy (nali):

sorry the cos C is supposed to be next to the rest of the equation

OpenStudy (nali):